Two cars travelling in opposite directions were 360 km apart at the end of two hrs. If one car travels 6kph faster than the other, find the average speed of the slower car.
Answer by mananth(12270)

speed of one car = x mph
speed of second car = x+6 mph
..
they are in opposite directions
so add up the speeds
x+x+6 = 2x+6 mph.
distance = 240 miles
..
r=d/t
r=240/(2x+6)
2=240/(2x+6)
multiply by (2x+60
2(2x+6)=240
4x+12=240
-12
4x=240-12
4x=228
/4
x=57 mph speed of one car
speed of other car = x+6 = 57+6 = 63 mph
..
m.ananth@hotmail.ca